# Aerial Grasping via Maximizing Delta-Arm Workspace Utilization

### v1

@@ Abstract
The workspace limits the operational capabilities and range of motion for the systems with robotic arms. Maximizing workspace utilization has the potential to provide more optimal solutions for aerial manipulation tasks, increasing the system’s flexibility and operational efficiency. In this paper, we introduce a novel planning framework for aerial grasping that maximizes workspace utilization. We formulate an optimization problem to optimize the aerial manipulator’s trajectory, incorporating task constraints to achieve efficient manipulation. To address the challenge of incorporating the delta arm’s non-convex workspace into optimization constraints, we leverage a Multilayer Perceptron (MLP) to map position points to feasibility probabilities. Furthermore, we employ Reversible Residual Networks (RevNet) to approximate the complex forward kinematics of the delta arm, utilizing efficient model gradients to eliminate workspace constraints. We validate our methods in simulations and real-world experiments to demonstrate their effectiveness.
@@ I Introduction
Previous research in aerial manipulators has primarily focused on developing precise control [4][5] and advancing hardware design [6][7] to expand application scopes and capability boundaries. However, the planning component in manipulation systems is of equal importance yet has received insufficient research attention. In this paper, we focus on grasp planning for aerial manipulators with delta arms, ensuring compliance with dynamic, kinematic, and task constraints (c.f. Fig. 1).
@@
Figure 2: The overview of our proposed manipulation planning framework for aerial manipulators with delta arms, which mainly consists of four modules. (a) The initialization utilizes a pathfinder to generate a reference trajectory traversing all target points. (b) The workspace representation module leverages a learning-based method to map coordinate points to feasibility possibilities. (c) The whole-body optimization problem is formulated by continuous polymathic trajectory generation and task constraint for manipulation tasks. (d) The workspace constraint elimination module approximates the forward kinematics of the delta arm with a learning model to calculate mapping and gradient. Finally, the aerial manipulator performs successful aerial grasping in real-world experiments.
@@ II-B Workspace Constraint Modeling
The robotic arm’s workspace plays a crucial role in trajectory optimization by preventing the inclusion of unreachable points [19]. Conventional feasible space constraint formulation methods are typically limited to convex polyhedral regions in three-dimensional position space, which simplifies implementation but restricts flexibility. This limitation becomes particularly problematic for delta arms, which possess inherently non-convex workspaces (c.f. Fig. 3 (a)), making direct integration of delta workspace constraints impractical using traditional approaches. To address this challenge, various approximation methods have been proposed. Deng et al. [8] approximate the workspace using a slender rectangular prism extending along the z-axis in the arm’s coordinate system. Although this approach preserves convexity and fully utilizes the manipulator’s motion capabilities along the $z$-axis, it fails to exploit the full range of motion along the $x$ and $y$ axes, limiting overall system performance. Similarly, Cao et al. [9] employ the maximum inscribed rectangular prism within the workspace, which provides a simple representation but still achieves relatively low workspace utilization. To overcome these limitations, our method utilizes deep learning to map 3D coordinates to feasibility probabilities, enabling direct optimization within the entire non-convex workspace and enhancing the system’s motion flexibility.
@@ III-A Coordinates Definition and Dynamic Model
$\begin{cases}\bm{f}=\ddot{\bm{p}}_{b}+g\bm{z}_{\mathcal{W}}\\ \bm{\tau}=\mathbf{I}\cdot{}^{B}\dot{\bm{\omega}}_{b}+{}^{B}\bm{\omega}_{b}\times\left(\mathbf{I}\cdot{}^{B}\bm{\omega}_{b}\right)\\ \dot{\mathbf{R}}_{b}=\mathbf{R}_{b}{}^{B}\hat{\bm{w}}_{b}\\ {}^{A}\bm{p}_{e}=h_{e}(\bm{\theta})\end{cases}$ (1)
where $\bm{f},\bm{\tau}\in\mathbb{R}^{3}$ represent the total normalized thrust and total torque of the quadrotor, respectively. $g$ is the gravitational acceleration, and $\bm{z}_{\mathcal{W}}$ is the z-axis of $\mathcal{F}_{W}$. $\mathbf{I}$ represents the moment of the inertia matrix, and ${}^{B}\bm{\omega}_{b}$ is the body angular velocity in $\mathcal{F}_{B}$. $\hat{\bm{w}}_{b}$ is the skew-symmetric matrix form of $\bm{w}_{b}$. $\bm{\theta}=[\theta_{1}\quad\theta_{2}\quad\theta_{3}]^{T}$ is the three joint angles of the delta arm ($\theta_{i}\in[0,90^{\circ}],i=1,2,3$), and $h_{e}(\cdot):\mathbb{R}^{3}\to\mathbb{R}^{3}$ denotes the forward kinematics.
@@ III-B Trajectory Representation
In this paper, we adopt the MINCO representation [21], using piece-wise polynomial trajectory to conduct spatial-temporal deformation. The entire trajectory for aerial manipulators with delta arms can be represented as:
@@
where $\mathbf{c}_{i}\in\mathbb{R}^{2s\times 6}$ is the coeffient matrix of the $i^{th}$ piece. For computational convenience, we split $\mathbf{c}_{i}$ into $\mathbf{c}_{b,i}\in\mathbb{R}^{2s\times 3}$ for the aerial manipulator and $\mathbf{c}_{e,i}\in\mathbb{R}^{2s\times 3}$ for the end-effector. $s$ represents the order of the integrator chain. $\mathbf{\beta}(t)=[1\quad t\quad t^{2}\quad\cdots\quad t^{2s-1}]^{T}$ is the natural basis vector and $T_{i}=t_{i}-t_{i-1}$ is the duration of the $i^{th}$ piece.
@@ IV-A Workspace Representation
Previous approaches employs an inscribed cube to approximate the workspace [9], providing an efficient representation for penalty and gradient calculation. However, this approximation significantly reduces the end-effector’s feasible space during planning, limiting the aerial manipulator’s manipulation flexibility. Furthermore, the exclusion of the non-convex portion (mainly the upper portion) of the workspace forces the robotic arm to maintain a highly extended state during flight, which compromises control accuracy and introduces potential risks. To overcome these limitations, we propose a deep learning-based method to represent the delta workspace, which enables maximum utilization of the entire workspace in the planning process. The comparison of the two methods is shown in Fig. 3 (a).
To map the relative positional relationship between ${}^{A}\bm{p}_{e}$ and the workspace, we model the workspace as:
@@
With the strong approximation capability of neural networks [22], to derive this complex nonlinear function $F_{w}$, we design a lightweight 6-layer MLP, as shown in Fig. 4 (a). The three-dimensional coordinates are input to the network and transformed through five fully connected layers: first expanding to 64 neurons, then to 256 neurons, followed by reduction to 128 neurons, further reduction to 64 neurons, and finally outputting a single value. LeakyReLU activation functions are employed between each fully connected layer, introducing nonlinear elements after each linear transformation, enabling the network to learn complex nonlinear mapping relationships. The final layer uses a Sigmoid activation function to compress the one-dimensional output value into the range of 0-1, conforming to the range constraint of probability measures.
@@
To quantitatively compare the workspace utilization efficiency between our method and that of [9], we implemented a testing program using MATLAB. We reconstruct the delta workspace using the alpha shape method, compute its volume, and compare it with the inscribed cube volume. The analysis reveals that the volume of the inscribed cube is 551 cm3, while our method achieves 2553 cm3, representing a 363% increase over the previous approach. This substantial improvement in workspace utilization enables the delta arm to execute more flexible movements during aerial manipulation tasks.
@@ IV-B1 Formulation
To ensure the smoothness, feasibility of the trajectory, and its effective adaptation to manipulation task requirements, we construct the objective function of trajectory optimization as followed:
@@ IV-B2 Task Constraints
$\mathcal{J}_{\mathcal{G}_{K},i}$ =$=\begin{cases}0,&i\notin\mathcal{I}\\ \mathcal{S}(\mathcal{G}_{p,i})+\mathcal{S}(\mathcal{G}_{o,i}),&i\in\mathcal{I}\end{cases}$ (11)
@@
where $\mathcal{I}$ denotes the set of specific trajectory segment indices designated for executing the grasp actions. $\mathcal{S}(\cdot)$ is the cubic smooth function given by $\mathcal{S}(x)=\text{max}(x^{3},0)$. $t_{\gamma}\in[0,T_{i}]$, which is the temporal point closest to the target waypoint. $e_{o}$ is the maximum allowable angle difference, and $\bm{f}_{i}\in\mathbb{R}^{3}$ is the normalized thrust vector.
@@
$\mathcal{J}_{\mathcal{G}_{K},i}=\begin{cases}0,&i\notin\mathcal{I}\\ \mathcal{S}(\|\dot{\bm{p}}_{e,i}(t_{\gamma})\|^{2}),&i\in\mathcal{I}\end{cases}.$ (14)
@@
Leveraging the gradient of the workspace representation model, we derive the gradient of workspace penalty with respect to the end-effector position in $\mathcal{F}_{A}$:
@@ IV-C Workspace Constraint Elimination
In our framework, all trajectories can be parameterized by time vector $\mathbf{T}$ and intermediate points $\mathbf{q}=[\bm{q}_{1}\cdots\bm{q}_{M-1}]^{T}$ where $\bm{q}_{i}=[\bm{q}_{b,i}^{T}\quad{}^{A}\bm{q}_{e,i}^{T}]^{T}$. These parameters are subject to specific constraints: $\bm{q}_{b,i}$ must lie within the polyhedrons, ${}^{A}\bm{q}_{e,i}$ must remain within the delta arm workspace, and all elements of the time vector $\mathbf{T}$ must be positive. To improve optimization efficiency when handling constrained variables, Wang et al. [21] proposed a polyhedron elimination method that transforms the constrained variables $\bm{q}_{b,i}$ and $\mathbf{T}$ into unconstrained variables. However, this method is applicable only when the constraint space for each point is convex, making it unsuitable for handling the non-convex workspace constraints of ${}^{A}\bm{q}_{e,i}$ for the delta arm. To address this limitation, we propose a deep learning-based workspace constraint elimination method specifically designed for the delta arm, which effectively transforms ${}^{A}\bm{q}_{e,i}$ into unconstrained variables while maintaining efficient gradient computation throughout the optimization process.
It is noteworthy that all the Cartesian coordinates within the delta arm’s workspace can be independently computed through forward kinematics by servo joint angles $\bm{\theta}$ (Equation 1). To transform these joint angles into unconstrained variables for optimization, we define two variables $[\delta_{n},\zeta_{n}]$ for each joint angle where $n\in{1,2,3}$. For convenience, we normalize each joint angle to the range of 0-1 using:
@@
During the optimization process, update $\mathbf{\Xi}$ through penalty requires gradient propagation, which is given by:
@@
Due to the relatively complex forward kinematics of the delta arm, direct calculations of $\frac{\partial{}^{A}\bm{q}_{e}}{\partial\bm{\theta}}$ may result in numerical instabilities, singularity issues, and computational inefficiency, making it difficult to apply in practice. To address this challenge, we propose using deep learning networks to approximate the forward kinematics, enabling stable gradient computation through the inherent differentiability of the network architecture.
To approximate the complex forward kinematics using the lightest possible network, we designed Delta RevNet (c.f. Fig. 4 (b)) with reference to RevNet [23]. The model takes a three-dimensional vector with values in the range [0,1], which is $\bm{\vartheta}=[\vartheta_{1}\quad\vartheta_{2}\quad\vartheta_{3}]^{T}$ as input and ${}^{A}\bm{q}_{e}$ as output. The core of our network is the RevBlock (reversible block) structure. Each RevBlock evenly divides the input tensor along the feature dimension into two parts, then performs nonlinear transformations through two sequential networks, F and G, each composed of fully connected layers, batch normalization, and ReLU activation functions.
@@ V-A Model Training
The models used in all experiments are trained on a server with an NVIDIA RTX4090 GPU and CUDA 12.4. For the first model (c.f. Fig. 4 (a)), we generate the dataset by randomly sampling points within a bounded three-dimensional space and determining whether each point lies within the workspace through inverse kinematics calculations. Points outside the workspace are labeled as 0, while those inside are labeled as 1. For the second model (c.f. Fig. 4 (b)), we sample 3D vectors within the range $[0,1]$, denormalize them by scaling to 90, and compute the end-effector position through forward kinematics as labels. The two models are trained for 50 and 100 epochs, respectively, achieving excellent performance on the test set with error rates below 0.001% across 100,000 test samples.

### v2

@@ Abstract
Workspace limitations restrict the operational capabilities and range of motion for systems with robotic arms. Maximizing workspace utilization has the potential to provide better solutions for aerial manipulation tasks, increasing the system’s flexibility and operational efficiency. In this paper, we introduce a novel planning framework for aerial grasping that maximizes workspace utilization. We formulate an optimization problem to optimize the aerial manipulator’s trajectory, incorporating task constraints to achieve efficient manipulation. To address the challenge of incorporating the delta arm’s non-convex workspace into optimization constraints, we leverage a Multilayer Perceptron (MLP) to map the point positions to feasibility probabilities. Furthermore, we employ Reversible Residual Networks (RevNet) to approximate the complex forward kinematics of the delta arm, utilizing its efficient model gradients to further eliminate workspace constraints. We validate our methods in simulations and real-world experiments to demonstrate their effectiveness.
@@ I Introduction
Previous research in aerial manipulators has primarily focused on developing precise control [4][5] and advancing hardware design [6][7] to expand application scopes and capability boundaries. However, the planning component in manipulation systems is of equal importance, yet has received insufficient research attention. In this paper, we focus on grasp planning for aerial manipulators with delta arms, ensuring compliance with dynamic, kinematic, and task constraints (cf. Fig. 1).
@@
Figure 2: The overview of our proposed manipulation planning framework for aerial manipulators with delta arms, which mainly consists of four modules. (a) The initialization utilizes a pathfinder to generate a reference path traversing all target points. (b) The workspace representation module leverages a learning-based method to map coordinate points to feasibility possibilities. (c) The whole-body optimization problem is formulated by continuous polymathic trajectory generation and task constraints for manipulation tasks. (d) The workspace constraint elimination module approximates the forward kinematics of the delta arm with a learning model to calculate mapping and gradient. Finally, the aerial manipulator performs successful aerial grasping in real-world experiments.
@@ II-B Workspace Constraint Modeling
The robotic arm’s workspace plays a crucial role in trajectory optimization by preventing the inclusion of unreachable points [19]. Conventional feasible space constraint formulation methods are typically limited to convex polyhedral regions in three-dimensional position space, which simplifies implementation but restricts flexibility. This limitation becomes particularly problematic for delta arms, which possess inherently non-convex workspaces (cf. Fig. 3 (a)), making direct integration of delta workspace constraints impractical using traditional approaches. To address this challenge, various approximation methods have been proposed. Deng et al. [8] approximate the workspace using a slender rectangular prism extending along the z-axis in the arm’s coordinate system. Although this approach preserves convexity and fully utilizes the manipulator’s motion capabilities along the $z$-axis, it fails to exploit the full range of motion along the $x$ and $y$ axes, limiting overall system performance. Similarly, Cao et al. [9] employ the maximum inscribed rectangular prism within the workspace, which provides a simple representation but still achieves relatively low workspace utilization. To overcome these limitations, our method utilizes a neural network to map 3D coordinates to feasibility probabilities, enabling direct optimization within the entire non-convex workspace and enhancing the system’s motion flexibility.
@@ III-A Coordinates Definition and Dynamic Model
$\begin{cases}\bm{f}=\ddot{\bm{p}}_{b}+g\bm{z}_{\mathcal{W}}\\ \bm{\tau}=\mathbf{I}\cdot{}^{B}\dot{\bm{\omega}}_{b}+{}^{B}\bm{\omega}_{b}\times\left(\mathbf{I}\cdot{}^{B}\bm{\omega}_{b}\right)\\ \dot{\mathbf{R}}_{b}=\mathbf{R}_{b}{}^{B}\hat{\bm{\omega}}_{b}\\ {}^{A}\bm{p}_{e}=h_{e}(\bm{\theta})\end{cases}$ (1)
where $\bm{f},\bm{\tau}\in\mathbb{R}^{3}$ represent the total normalized thrust and total torque of the quadrotor, respectively. $g$ is the gravitational acceleration, and $\bm{z}_{\mathcal{W}}$ is the z-axis of $\mathcal{F}_{W}$. $\mathbf{I}$ represents the moment of the inertia matrix, and ${}^{B}\bm{\omega}_{b}$ is the body angular velocity in $\mathcal{F}_{B}$. ${}^{B}\hat{\bm{\omega}}_{b}$ is the skew-symmetric matrix form of ${}^{B}\bm{\omega}_{b}$. $\bm{\theta}=[\theta_{1}\quad\theta_{2}\quad\theta_{3}]^{T}$ is the three joint angles of the delta arm ($\theta_{i}\in[0,90],i=1,2,3$), and $h_{e}(\cdot):\mathbb{R}^{3}\to\mathbb{R}^{3}$ denotes the forward kinematics.
@@ III-B Trajectory Representation
In this paper, we adopt the MINCO representation [21], using a piece-wise polynomial trajectory to conduct spatial-temporal deformation. The entire trajectory for aerial manipulators with delta arms can be represented as:
@@
where $\mathbf{c}_{i}\in\mathbb{R}^{2s\times 6}$ is the coefficient matrix of the $i^{th}$ piece. For computational convenience, we split $\mathbf{c}_{i}$ into $\mathbf{c}_{b,i}\in\mathbb{R}^{2s\times 3}$ for the aerial manipulator and $\mathbf{c}_{e,i}\in\mathbb{R}^{2s\times 3}$ for the end-effector. $s$ represents the order of the integrator chain. $\mathbf{\beta}(t)=[1\quad t\quad t^{2}\quad\cdots\quad t^{2s-1}]^{T}$ is the natural basis vector and $T_{i}=t_{i}-t_{i-1}$ is the duration of the $i^{th}$ piece.
@@ IV-A Workspace Representation
Previous approaches employ an inscribed cube to approximate the workspace [9], providing an efficient representation for penalty and gradient calculation. However, this approximation significantly reduces the end-effector’s feasible space during planning, limiting the aerial manipulator’s manipulation flexibility. Furthermore, the exclusion of the non-convex portion (mainly the upper portion) of the workspace forces the robotic arm to maintain a highly extended state during flight, which compromises control accuracy and introduces potential risks. To overcome these limitations, we propose a learning-based method to represent the delta workspace, which enables maximum utilization of the entire workspace in the planning process. The comparison of the two methods is shown in Fig. 3 (a).
The optimization process requires determining the relative position between ${}^{A}\bm{p}_{e}$ and the workspace, which can be approximated as a binary classification problem using feasibility probability to distinguish points inside and outside the workspace. The problem can be modeled as:
@@
With the strong approximation capability of neural networks [22], to derive this complex nonlinear function $F_{w}$, we design a lightweight 6-layer MLP, as shown in Fig. 4 (a). The three-dimensional coordinates are input to the network and transformed through five fully connected layers: first expanding to 64 neurons, then to 256 neurons, followed by reduction to 128 neurons, further reduction to 64 neurons, and finally outputting a single value. ReLU activation functions are employed between each fully connected layer, introducing nonlinear elements after each linear transformation, enabling the network to learn complex nonlinear mapping relationships. The final layer uses a Sigmoid activation function to compress the one-dimensional output value into the range of [0,1][0,1][ 0 , 1 ], conforming to the range constraint of probability measures.
@@
To quantitatively compare the workspace utilization efficiency between our method and that of [9], we implemented a testing program using MATLAB. We reconstruct the delta workspace using the alpha shape method with $\alpha=10$, compute its volume, and compare it with the inscribed cube volume. The analysis reveals that the volume of the inscribed cube is 551 cm3, while our method achieves 2553 cm3, representing a 363% increase over the previous approach. This substantial improvement in workspace utilization enables the delta arm to execute more flexible movements during aerial manipulation tasks.
@@ IV-B1 Formulation
To ensure the smoothness, feasibility of the trajectory, and its effective adaptation to manipulation task requirements, we construct the objective function of trajectory optimization as follows:
@@ IV-B2 Task Constraints
$\mathcal{J}_{\mathcal{G}_{K},i}$ =$=\begin{cases}\mathcal{S}(\mathcal{G}_{p,i})+\mathcal{S}(\mathcal{G}_{o,i}),&t=t_{\gamma},\\ 0,&\text{otherwise}.\\ \end{cases}$ (11)
@@
where $t_{\gamma}=\sum^{l-1}_{j=0}T_{j}$ is the instant closest to the specific waypoint, $l\in\mathcal{I}$ denotes the index of the trajectory segment designated for executing the grasp actions, and $\mathcal{I}$ denotes the set of such segment indices. $\mathcal{S}(\cdot)$ is the cubic smooth function given by $\mathcal{S}(x)=\text{max}(x^{3},0)$. $e_{o}$ is the maximum allowable angle difference, and $\bm{f}_{i}\in\mathbb{R}^{3}$ is the normalized thrust vector.
@@
$\mathcal{J}_{\mathcal{G}_{K},i}=\begin{cases}0,&t=t_{\gamma},\\ \mathcal{S}(\|\dot{\bm{p}}_{e,i}(t_{\gamma})\|^{2}),&\text{otherwise}.\end{cases}$ (14)
@@
Leveraging the gradient of the workspace representation model, we derive the gradient of the workspace penalty with respect to the end-effector position in $\mathcal{F}_{A}$:
@@ IV-C Workspace Constraint Elimination
In our framework, all trajectories can be parameterized by the time vector $\mathbf{T}$ and intermediate points $\mathbf{q}=[\bm{q}_{1}\cdots\bm{q}_{M-1}]^{T}$ where $\bm{q}_{i}=[\bm{q}_{b,i}^{T}\quad{}^{A}\bm{q}_{e,i}^{T}]^{T}$. These parameters are subject to specific constraints: $\bm{q}_{b,i}$ must lie within the polyhedrons, ${}^{A}\bm{q}_{e,i}$ must remain within the delta arm workspace, and all elements of the time vector $\mathbf{T}$ must be positive. To handle the hard constraints on optimization variables, Wang et al. [21] proposed a polyhedron elimination method that transforms the constrained variables $\bm{q}_{b,i}$ and $\mathbf{T}$ into unconstrained variables. However, this method is applicable only when the constraint space for each point is convex, making it unsuitable for handling the non-convex workspace constraints of ${}^{A}\bm{q}_{e,i}$ for the delta arm. To address this limitation, we propose a learning-based workspace constraint elimination method specifically designed for the delta arm, which effectively transforms ${}^{A}\bm{q}_{e,i}$ into unconstrained variables while maintaining efficient gradient computation throughout the optimization process.
It is noteworthy that all the Cartesian coordinates within the delta arm’s workspace can be independently computed through forward kinematics by servo joint angles $\bm{\theta}$ (Equation 1). To transform these joint angles into unconstrained variables for optimization, we define two variables $[\delta_{n},\zeta_{n}]$ for each joint angle where $n\in\{1,2,3\}$. For convenience, we normalize each joint angle to the range of [0,1][0,1][ 0 , 1 ] using:
@@
During the optimization process, updating $\mathbf{\Xi}$ through penalty terms requires gradient propagation, which is given by:
@@
Due to the relatively complex forward kinematics of the delta arm, direct calculations of $\frac{\partial{}^{A}\bm{q}_{e}}{\partial\bm{\theta}}$ may lead to numerical instabilities, especially when the inverse kinematics is infeasible, thereby limiting its practical applicability. To address this challenge, we propose using neural networks to approximate the forward kinematics, enabling stable gradient computation through the inherent differentiability of the network architecture.
To approximate the complex forward kinematics using the lightest possible network, we designed Delta RevNet (cf. Fig. 4 (b)) with reference to RevNet [23]. The model takes a three-dimensional vector with values in the range [0,1], which is $\bm{\vartheta}=[\vartheta_{1}\quad\vartheta_{2}\quad\vartheta_{3}]^{T}$ as input and ${}^{A}\bm{q}_{e}$ as output. The core of our network is the RevBlock (reversible block) structure. Each RevBlock evenly divides the input tensor along the feature dimension into two parts, then performs nonlinear transformations through two sequential networks, F and G, each composed of fully connected layers, batch normalization, and ReLU activation functions.
@@
Finally, the simplified optimization problem is addressed through the L-BFGS method.
@@ V-A Model Training
The models used in all experiments are trained on a server with an NVIDIA RTX4090 GPU and CUDA 12.4. For the first model (cf. Fig. 4 (a)), we generate the dataset by randomly sampling points within a bounded three-dimensional space and determining whether each point lies within the workspace through inverse kinematics calculations. Points outside the workspace are labeled as 0, while those inside are labeled as 1. For the second model (cf. Fig. 4 (b)), we sample 3D vectors within the range [0,1][0,1][ 0 , 1 ], denormalize them by scaling to 90, and compute the end-effector position through forward kinematics as labels. The two models are trained for 50 and 100 epochs, respectively, achieving excellent performance on the test set with error rates below 0.001% across 100,000 test samples.